Bromine monochloride, BrCl decomposes into bromine and chlorine and reaches the equilibrium: 2BrCl (g) = Br2 (g) + Cl2 (g) for which Kc= 32 at 500 K. If initially pure BrCl is present at a concentration of 3.30 mbar, what is its partial pressure in the mixture at equilibrium

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Already ANswered this:
Plz check your previous question,
 
            2BrCl (g) = Br2 (g) + Cl2 (g) 
t=0            3.3 * 10^-3    0              0
t=t           3.3 * 10^-3 – x   x/2            x/2
Kc = x^2/4(3.3 * 10^-3 – x)^2 = 32
x/(3.3 * 10^-3 – x) = 2 x 4 x root(2)
x =  37.3 –  11.3x
12.3x = 37.3
x = 3.035 mBar.
At eqm. initial partial pressure would be hence,
x/2 for each Br2 and Cl2 that is = > 1.517 mBar.